# Invariant derivations and trace bounds

### v3

@@ 7. Corners and the rank of $M$
Continue the assumption of sections 5 and 6 that $G$ is a connected reductive algebraic group whose derived subgroup is simply connected. We now investigate the significance of the rank of $M$ at an element $g$. Note that $M$ and $M^{\sigma}$ have the same rank at $g$, because $M^{\sigma}=J(\sigma)^{\mathsf{T}}M$, where $\sigma$ is an automorphism, so $J(\sigma)$ is invertible.
@@
$\Omega_{\mathbb{C}[T/W]}\otimes_{\mathbb{C}[T/W]}\mathbb{C}[T]\to\Omega_{\mathbb{C}[T]}$
has image $(\Omega_{\mathbb{C}[T]}|_{t})^{W_{t}}$.
Since $T$ is abelian, the map $t\mapsto w(t)t^{-1}$ for $w\in W$ is a homomorphism $T\mapsto T$. In the proof, we will use the notation $H_{0}(W;T)$ for the coinvariants, i.e., $T$ modulo the subgroup consisting of elements $w(t)t^{-1}$ for $w\in W$ and $t\in T(\mathbb{C})$ [Wei94, Ch. 6]. Note that the quotient $H_{0}(W;T)$ of $T$ is itself an algebraic group over $\mathbb{C}$, and therefore is itself a torus.
@@
$\Omega_{\mathbb{C}[T/W]}\otimes_{\mathbb{C}[T/W]}\mathbb{C}[T]\to\Omega_{\mathbb{C}[T/W_{t}]}\otimes_{\mathbb{C}[T/W_{t}]}\mathbb{C}[T]\to\Omega_{\mathbb{C}[T]}.$
Since $T/W_{t}\to T/W$ is étale at $t$, the first map is an isomorphism near $t$, and thus we may assume without loss of generality that $W_{t}=W$, i.e., that $t$ is fixed by $W$. Moreover, since the map is equivariant under translations by $T^{W}$, we may assume that $t=1$.
In this case, the map is $W$-equivariant, and thus the restriction to $1$ is still $W$-equivariant; since $W$ acts trivially on $\Omega_{\mathbb{C}[T/W]}$, it follows that the image is contained in $(\Omega_{\mathbb{C}[T]}|_{1})^{W}$. That space is the same as $(\mathfrak{t}^{*})^{W}$ since $\Omega_{\mathbb{C}[T]}|_{1}$ is canonically identified with $\mathfrak{t}^{*}$.
To get the other inclusion, consider the natural map $T\to H_{0}(W;T)$. This again induces a pullback map on differentials. The fiber at $1$ of the map
$\Omega_{\mathbb{C}[H_{0}(W;T)]}\otimes_{\mathbb{C}[H_{0}(W;T)]}\mathbb{C}[T]\to\Omega_{\mathbb{C}[T]}$
is the base change of the natural injective map
@@
which has image $(\mathfrak{t}^{*})^{W}$. Since the map $T\to H_{0}(W;T)$ factors through $T/W$, the map on differentials factors through the map of interest, and thus its image contains $(\mathfrak{t}^{*})^{W}$ as required. ∎
@@
For any $t\in T(\mathbb{C})$, the rank of $M(t)$ is equal to the dimension of $H_{0}(W_{t};T)$.
@@
The rank of $M(t)$ is the same as the rank of the restriction of $A$ to the image of the fiber at $t$ of
$\Omega_{\mathbb{C}[T/W]}\otimes_{\mathbb{C}[T/W]}\mathbb{C}[T]\to\Omega_{\mathbb{C}[T]},$
which we have just seen is nothing other than $(\mathfrak{t}^{*})^{W_{t}}$. Since $A$ is nonsingular and $W_{t}$-invariant, its restriction to the subspace of $W_{t}$-invariants is again nonsingular, so has rank
$\dim((\mathfrak{t}^{*})^{W_{t}})=\dim(H_{0}(W_{t};\mathfrak{t}))=\dim(H_{0}(W_{t};T)).\qed$
Proposition 7.4.
The rank of $M$ at a semisimple $g\in G(\mathbb{C})$ is equal to the dimension of the center of the centralizer of $g$.
There is an analogue of this result in the setting of Lemma 6.9. A linear version of 7.4 is part of the main theorem of [Ste60].
Proof.
We may reduce to the case of a (complex) point of the maximal torus. The claim to be proved is equivalent to saying that the corank of $M$ equals the rank of the semisimple part of $Z_{G}(g)$. This latter rank equals the rank of the stabilizer of $g$ in the Weyl group (compare, for example, [Ree10, §2.4]), which is itself complementary to the dimension of the space of $W_{g}$-invariants in the Cartan subalgebra. Since
$\dim(\mathfrak{t}^{W_{g}})=\dim(H_{0}(W_{g};T)),$
the claim follows. ∎
@@
By hypothesis, $G$ is reductive, so the identity component $Z_{G}(G)^{\circ}$ of the center of $G$ is a torus $T$. For every $g\in G(\mathbb{C})$, the center of $Z_{G}(g)$ contains $T$. On the other hand, if $g$ is a corner, then by the proposition $\dim Z(Z_{G}(g))=0$, so $T=1$. ∎
For context, we mention the following result, which provides another view on the preceding corollary and Cor. 5.12:
@@
Corners and simple groups
In view of Corollary 7.5, we now suppose that $G$ is simple compact (and, continuing from before, simply connected) with maximal torus $T$. Then $T/W$ is homeomorphic to the closure of an alcove in the coroot space [Bou05, §IX.5.2, Cor. 1]. The reasoning in the proof of Prop. 7.4 shows that the conjugacy classes of corners in $G$ correspond to vertices of the closure of the alcove, hence the name “corners”. The vertices consist of the element 0 (corresponding to the class of the identity in $T/W$) or an element of the form $a^{-1}\omega^{\vee}$, where $\omega^{\vee}$ is a fundamental coweight and $a$, an integer, is the coefficient of the simple root corresponding to $\omega^{\vee}$ in the expression of the highest root as a sum of simple roots [Ree10, §2.2]. In particular, we find:
@@
If $G$ is simple simply connected compact of rank $r$, then it has $r+1$ conjugacy classes of corners and each consists of elements of finite order. For every $g$ in the conjugacy class of a corner, so is $g^{m}$ for every $m$.
@@
The paragraph preceding the statement showed the first claim. So suppose that $g$ corresponds to an element $a^{-1}\omega^{\vee}$. Then $g^{a}$ is central in $G$ (by applying [Bou05, §IX.5.2, Cor. 1] for the adjoint group $G/Z(G)$), so it follows that the order of $g$ divides $a$ times the exponent of $Z(G)$. The centralizer of $g$ is a semisimple subgroup of $G$ of rank $r$ as in the proof of Prop. 7.4. It follows that the centralizer of $g^{m}$, which contains $Z_{G}(g)$, is also semisimple (see [Leh16, §6] for a discussion of such centralizers), so $g^{m}$ is also a corner. ∎
@@
Suppose $G$ is simple simply connected. Then $-1$ is in the Weyl group if and only if every character takes integer values on the corners.
@@
Now, there is nothing to prove if $g$ is the identity element, so we may assume that $g$ corresponds to a vertex $a^{-1}\omega^{\vee}$ in the alcove and therefore $m$ divides $a$ times the exponent of the center of $G$. This disposes of the cases where $G$ has type $G_{2}$ or $F_{4}$. For types $B$, $C$, and $D_{2n}$, $a$ is 1 or 2 while central elements have order dividing 2, so $m$ divides 4.
For $E_{7}$ and $E_{8}$, the same argument treats all the corners except one in each case, the one where $\omega^{\vee}$ is dual to the unique simple root in the Dynkin diagram that has three neighbors. However, the Galois group of $\mathbb{Q}(\cos(2\pi/m))$ permutes the character values $\chi(g),\chi(g^{2}),\ldots$. Since every power of $g$ is a corner (Cor. 7.7) and $\chi$ takes integers values on all the other corners, $\chi(g)$ must be fixed by the Galois group and so belongs to $\mathbb{Q}$.
@@ 8. Fundamental representations of $F_{4}$ and $E_{8}$
In Examples 5.18 and 7.8, we applied the method of invariant derivations to determine the range of trace values on the fundamental irreducible representations of the compact simple Lie group of type $G_{2}$. In this section, we will do the same for $F_{4}$ and $E_{8}$, although for $E_{8}$ we will only give complete information on the fundamental representations of dimension less than 100 million.
@@
Step 1: Compute the value of $f_{i}$ on the corners, i.e., the vanishing locus of $M$. There are $r+1$ corners for $G$ of rank $r$ (Cor. 7.7). For $F_{4}$ and $E_{8}$, we list the results in Tables 3 and 4, where the header row of the table gives the value of $i$. In the table, we have sorted the rows so that the first corner, i.e., the first data row, gives the dimension $d:=f_{i}(1)$ of the representation. In column $i$, we underline the minimum value $m$ of $f_{i}$ on a corner.
@@ 9. Simple groups of exceptional type
To find the critical points, we use that critical points of $\operatorname{Tr}\operatorname{Ad}$ is where the Weyl-invariant differentials vanish, which by Example 5.1 are the common solutions of
@@
The points of this scheme are the critical points in the complexification of a maximal torus. Some of these critical points have trace that lie outside the region claimed in Theorem 2.3, but that is no contradiction as long as they do not come from the compact group. A necessary condition for a toral element to come from the compact real group is that its trace on the adjoint representation of $A_{1}$ is inside the bound, i.e., $|t_{i}^{2}-2|\leq 2$. We look at all the critical points that satisfy this necessary condition and find that $\operatorname{Tr}\operatorname{Ad}$ is in the claimed range for all of them.
@@
There is a unique orbit of reflection hypertori, which are naturally maximal tori of subgroups of type $A_{5}$. This occurs as a component of a subgroup $A_{5}A_{1}$, and the corresponding decomposition of $\mathfrak{e}_{6}$ is:
@@
The nonregular elements of $A_{5}$ are given by a double cover of $\operatorname{GL}_{4}$ obtained by adjoining a square root $f_{2}$ of the determinant. We implicitly define functions $a_{0}$, $a_{1}$, and $b_{1}$ by writing the characteristic polynomial of the element of $\operatorname{GL}_{4}$ as
@@
The next case is isogenous to $A_{1}\times\mathbb{G}_{\mathrm{m}}^{2}$. We can then restrict the objective function to the maximal torus and simply solve for the vanishing of the gradient.

### v4

@@ 7. Corners and the rank of $M$
Continue the assumption of sections 5 and 6 that $G$ is a connected reductive algebraic group over $\mathbb{C}$ whose derived subgroup is simply connected. We now investigate the significance of the rank of $M$ at an element $g$. Note that $M$ and $M^{\sigma}$ have the same rank at $g$, because $M^{\sigma}=J(\sigma)^{\mathsf{T}}M$, where $\sigma$ is an automorphism, so $J(\sigma)$ is invertible.
@@
$\psi\colon\Omega_{\mathbb{C}[T/W]}\otimes_{\mathbb{C}[T/W]}\mathbb{C}[T]\to\Omega_{\mathbb{C}[T]}$
has image $(\Omega_{\mathbb{C}[T]}|_{t})^{W_{t}}$, which is canonically identified with $(\mathfrak{t}^{*})^{W_{t}}$.
Since $T$ is abelian, the map $t\mapsto w(t)t^{-1}$ for $w\in W$ is a homomorphism $T\mapsto T$. In the proof, we will use the notation $H_{0}(W;T)$ for the coinvariants, i.e., $T$ modulo the subgroup consisting of elements $w(t)t^{-1}$ for $w\in W$ and $t\in T(\mathbb{C})$ [Wei94, Ch. 6]. Note that the quotient $H_{0}(W;T)$ of $T$ is itself an algebraic group over $\mathbb{C}$, and therefore is itself a torus. The vector space $H_{0}(W;\mathfrak{t})$ is defined similarly.
@@
$\Omega_{\mathbb{C}[T/W]}\otimes_{\mathbb{C}[T/W]}\mathbb{C}[T]\to\Omega_{\mathbb{C}[T/W_{t}]}\otimes_{\mathbb{C}[T/W_{t}]}\mathbb{C}[T]\to\Omega_{\mathbb{C}[T]}$
of $\psi$. Since $T/W_{t}\to T/W$ is étale at $t$, the first map is an isomorphism near $t$, and thus we may assume without loss of generality that $W_{t}=W$, i.e., that $t$ is fixed by $W$. Moreover, since $\psi$ is equivariant under translations by $T^{W}$, we may assume that $t=1$.
In this case, $\psi$ is $W$-equivariant, and thus the restriction to $1$ is still $W$-equivariant; since $W$ acts trivially on $\Omega_{\mathbb{C}[T/W]}$, it follows that the image is contained in $(\Omega_{\mathbb{C}[T]}|_{1})^{W}$. That space is the same as $(\mathfrak{t}^{*})^{W}$ since $\Omega_{\mathbb{C}[T]}|_{1}$ is canonically identified with $\mathfrak{t}^{*}$.
To get the other inclusion, consider the natural quotient morphism $T\to H_{0}(W;T)$ of group schemes. This again induces a pullback map on differentials:
$\Omega_{\mathbb{C}[H_{0}(W;T)]}\otimes_{\mathbb{C}[H_{0}(W;T)]}\mathbb{C}[T]\to\Omega_{\mathbb{C}[T]}.$
The fiber at 1 of the pullback map is the dual of the corresponding morphism of Lie algebras,
@@
which is injective with image $({\mathfrak{t}}^{*})^{W}$. Since the morphism $T\to H_{0}(W;T)$ factors through the quotient scheme $T/W$, the map on differentials factors through $\psi$, and thus we have a factorization
$H_{0}(W;{\mathfrak{t}})^{*}\to\Omega_{\mathbb{C}[T/W]}|_{1}\to{\mathfrak{t}}^{*},$
so that the image of $\psi|_{1}$ contains the image $({\mathfrak{t}}^{*})^{W}$ of the composite. ∎
@@
For each semisimple $g\in G(\mathbb{C})$, the rank of $M(g)$ is equal to the dimension of the center of the centralizer of $g$.
There is an analogue of this result in the setting of Lemma 6.9. A linear version of 7.3 is part of the main theorem of [Ste60].
@@
We may reduce to the case of a (complex) point of the maximal torus. We show that each of the numbers in the statement is equal to the dimension of $(\mathfrak{t}^{*})^{W_{g}}$.
First consider the rank of $M(g)$. It is the same as the rank of the restriction of $A$ to the image of the fiber at $t$ of the map $\psi$ from Lemma 7.2, which we have just seen is nothing other than $(\mathfrak{t}^{*})^{W_{g}}$. Since $A$ is nonsingular and $W_{g}$-invariant, its restriction to the subspace of $W_{g}$-invariants is again nonsingular, proving the desired equality for the rank of $M(g)$.
For the other dimension, we use that the torus $T$ is also a maximal torus in $Z_{G}(g)$, so the stabilizer $W_{g}$ of $g$ is the Weyl group $N_{Z_{G}(g)}(T)/T$ of $Z_{G}(g)$. Now $Z_{G}(g)$ is reductive [Bor91, Prop. 13.19], so we may write $T=ZT^{\prime}$ where $Z$ is the identity component of the center of $Z_{G}(g)$ and $T^{\prime}$ is a maximal torus of the semisimple part of $Z_{G}(g)$. Each of these tori is stable under $W_{g}$, so
$\mathfrak{t}^{W_{g}}=\operatorname{Lie}(Z)^{W_{g}}\oplus\operatorname{Lie}(T^{\prime})^{W_{g}}=\operatorname{Lie}(Z),$
where $\operatorname{Lie}(T^{\prime})^{W_{g}}=0$ because the Weyl group of the semisimple group $Z_{G}(g)$ over $\mathbb{C}$ acts without fixed lines on the Lie algebra on the vector space generated by the coroots [Bou02, VI.1.2, Cor.].
Since $\mathfrak{t}$ has a non-degenerate $G$-invariant (hence also $W_{g}$-invariant) bilinear form, as representations of $W_{g}$, $\mathfrak{t}^{*}$ is equivalent to $\mathfrak{t}$, and so we have proved
$\dim\,(\mathfrak{t}^{*})^{W_{g}}=\dim\mathfrak{t}^{W_{g}}=\dim\operatorname{Lie}(Z),$
as required. ∎
@@
By hypothesis, $G$ is reductive, so the identity component $Z_{G}(G)^{\circ}$ of the center of $G$ is a torus $S$. For every $g\in G(\mathbb{C})$, the center of $Z_{G}(g)$ contains $S$. On the other hand, if the conjugacy class of $g$ is a corner, then by Prop. 7.3 $\dim Z(Z_{G}(g))=0$, so $S=1$. ∎
For context, we momentarily leave the domain of algebraic groups over $\mathbb{C}$ to mention the following result, which provides another view on the preceding corollary and Cor. 5.12:
@@
Return now to the setting where $G$ is an algebraic group over $\mathbb{C}$, and we seek to say more about the corners of $G$.
@@
Let $G$ be a semisimple and simply connected algebraic group over $\mathbb{C}$, and suppose $g\in G(\mathbb{C})$ is semisimple and its conjugacy class is a corner.
  1. (1)
    $g$ has finite order and the conjugacy class of $g^{m}$ is also a corner for every integer $m$.
  2. (2)
    If $G$ is simple of rank $r$, then $G$ has $r+1$ corners and the order of $g$ divides $a$ times the exponent of $Z(G)$, where $a$ is the coefficient of some simple root in the expression for the highest root of $G$ as a linear combination of simple roots.
@@
Put $H_{m}:=Z_{G}(g^{m})^{\circ}$. It is a reductive subgroup of $G$ and contains a maximal torus $T$ containing $g$. Evidently,
(7.7) $H_{m}\supseteq H_{1}\supseteq T.$
For every $m$, the center of the reductive group $H_{m}$ is contained in every maximal torus of $H_{m}$ [Mil17, Prop. 21.7], so we find
$T\supseteq Z(H_{1})\supseteq Z(H_{m}).$
Prop. 7.3 says that $H_{1}$ is semisimple, so $Z(H_{1})$ is finite, ergo $g$ has finite order. The reductive group $H_{m}$ also has finite center, so $H_{m}$ is semisimple (instead of merely reductive) and the class of $g^{m}$ is a corner by Prop. 7.3. This proves 1.
Now suppose that $G$ is simple of rank $r$. Since $g$ has finite order, it is conjugate to an element of a maximal compact subtorus $S$ of $T$. That subtorus is in turn homeomorphic to the closure of an alcove in the coroot space $V$ [Bou05, §IX.5.2, Cor. 1]. (The coroot space is the real vector space with basis the coroots dual to the simple roots of $G$ with respect to $T$.) Because $Z_{G}(g)^{\circ}$ is semisimple, its root system also has rank $r$, which means that the element $v\in V$ corresponding to $g$ is a vertex of the closure of the alcove, compare [Ree10, p. 11]. The vertices are of the following form. Put $\omega^{\vee}_{i}$ for the fundamental dominant weight of the dual root system, so that ${\left\langle{\omega^{\vee}_{i},\alpha_{j}}\right\rangle}=\delta_{ij}$ (Kronecker delta) for $\alpha_{1},\ldots,\alpha_{r}$ the simple roots of $G$ in a pinning involving $T$. Then $v$ is a vertex if and only if $v=0$ (in which case $g=1$) or $v=a_{i}^{-1}\omega^{\vee}_{i}$, where $a_{i}$ is the coefficient of the simple root $\alpha_{i}$ in the expression $\sum a_{j}\alpha_{j}$ for the highest root. In the second case, the corresponding $g$ has $g^{a_{i}}\in Z(G)$. This proves 2. ∎
Remark 7.8 (naming the corners).
Since a corner is a conjugacy classes of elements of finite order, when $G$ is simple the corner can be identified uniquely via its Kac coordinates as described in [Kac69], [Ser06], [Ree10, §2.2], or [Leh16, §6]. In the notation of the corollary, this is a list $(s_{0},s_{1},\ldots,s_{r})$ of non-negative integers, not all zero, with greatest common divisor 1. The corners are those conjugacy classes with Kac coordinates that have a 1 in one entry and zeros elsewhere; the corner $a_{i}^{-1}\omega^{\vee}_{i}$ has $s_{i}=1$.
Alternatively, suppose $G$ is simple of type $G_{2}$, $F_{4}$, or $E_{8}$. The conjugacy class of some $g\in G(\mathbb{C})$ is a corner and $p$ is a prime not dividing the order of $g$. One could use the methods of Appendix A to construct a finite-field analogue $G(\mathbb{F}_{p})$ of $G$ over the finite field $\mathbb{F}_{p}$ and relate $g$ with an element $\bar{g}$ of the same order in $G(\mathbb{F}_{p})$. In this way, it may make sense to refer to the conjugacy class of $g$ itself using a label 2A, 3B, etc., that is used for the conjugacy class of $\bar{g}$ in the style of the Atlas [CCN+85]. We do not pursue this here.
@@
Suppose $G$ is simple and simply connected. Then $-1$ is in the Weyl group if and only if every character takes integer values on the corners.
@@
By Corollary 7.62, $m$ divides $a$ times the exponent of center of $G$, for some coefficient $a$ appearing in the expression of the highest root. For types $G_{2}$ and $F_{4}$, the center is trivial and $a$ is 4 or divides 6, which disposes of those cases. For types $B$, $C$, and $D_{2n}$, $a$ is 1 or 2 while central elements have order dividing 2, so $m$ divides 4.
For $E_{7}$ and $E_{8}$, the same argument treats all the corners except one in each case, the one where $\omega^{\vee}$ is dual to the unique simple root in the Dynkin diagram that has three neighbors. However, the Galois group of $\mathbb{Q}(\cos(2\pi/m))$ permutes the character values $\chi(g),\chi(g^{2}),\ldots$. Since every power of $g$ is a corner (Cor. 7.61) and $\chi$ takes integer values on all the other corners, $\chi(g)$ must be fixed by the Galois group and so belongs to $\mathbb{Q}$.
@@ 8. Fundamental representations of $F_{4}$ and $E_{8}$
In Examples 5.20 and 7.9, we applied the method of invariant derivations to determine the range of trace values on the fundamental irreducible representations of the compact simple Lie group of type $G_{2}$. In this section, we will do the same for $F_{4}$ and $E_{8}$, although for $E_{8}$ we will only give complete information on the fundamental representations of dimension less than 100 million.
@@
Step 1: Compute the value of $f_{i}$ on the corners, i.e., the vanishing locus of $M$. There are $r+1$ corners for $G$ of rank $r$ (Cor. 7.6). For $F_{4}$ and $E_{8}$, we list the results in Tables 3 and 4, where the header row of the table gives the value of $i$. In the table, we have sorted the rows so that the first corner, i.e., the first data row, gives the dimension $d:=f_{i}(1)$ of the representation. In column $i$, we underline the minimum value $m$ of $f_{i}$ on a corner.
@@ 9. Simple groups of exceptional type
To find the critical points, we use that the critical points of $\operatorname{Tr}\operatorname{Ad}$ are where the Weyl-invariant differentials vanish, which by Example 5.1 are the common solutions of
@@
The points of this scheme are the critical points in the complexification of a maximal torus. Some of these critical points have trace that lie outside the region claimed in Theorem 2.3, but that is no contradiction as long as they do not come from the compact group. A necessary condition for a toral element to come from the compact group is that its trace on the adjoint representation of $A_{1}$ is inside the bound, i.e., $|t_{i}^{2}-2|\leq 2$. We look at all the critical points that satisfy this necessary condition and find that $\operatorname{Tr}\operatorname{Ad}$ is in the claimed range for all of them.
@@
There is a unique orbit of codimension 1 tori fixed pointwise by a reflection. These tori are naturally maximal tori of subgroups of type $A_{5}$. This occurs as a component of a subgroup $A_{5}A_{1}$, and the corresponding decomposition of $\mathfrak{e}_{6}$ is:
@@
Any nonregular semisimple element of $A_{5}$ is contained in a subgroup isomorphic to the double cover of $\operatorname{GL}_{4}$ obtained by adjoining a square root $f_{2}$ of the determinant (where the repeated eigenvalue is $f_{2}^{-1}$). We implicitly define functions $a_{0}$, $a_{1}$, and $b_{1}$ by writing the characteristic polynomial of the element of $\operatorname{GL}_{4}$ as
@@
The next case is isogenous to $A_{1}\times\mathbb{G}_{\mathrm{m}}^{2}$. We can then restrict $\operatorname{Tr}\operatorname{Ad}$ to the maximal torus in $A_{1}\times\mathbb{G}_{\mathrm{m}}^{2}$ and simply solve for the vanishing of the gradient.